OpenStudy (anonymous):

Evaluate the integral integrate (1+4cot x)/(4-cot x)dx

OpenStudy (jamesj):

Remember that cot x = cos x / sin x. Multiply numerator and denominator by sin x and you have \[\frac{ \sin x + 4 \cos x}{4 \sin x - \cos x}\] Now look carefully at these two terms and the answer will be apparent.